OpenStudy (anonymous):

2 over 3 (x - 2) = 4x Answers - 2 over 5 - 5 over 2 2 over 5 5

OpenStudy (the_fizicx99):

A

OpenStudy (lncognlto):

(2/3)(x - 2) = 4x (2/3)x - 4/3 = 4x 2x - 4 = 12x 10x = -4 x = -4/10 x = ?